### v1

@@ 5 Discussion
Vision-based fault diagnosis and self-recovery offered an effective means of enhancing the stability of strawberry-harvesting robots. During the inflating and approaching stage, the positional relationship between the gripper and the harvesting point served as a visual indicator of accumulated errors caused by fruit recognition, localization, hand–eye calibration, and inverse kinematics. Without requiring an additional camera, relative errors based on the simultaneous target-gripper detection were computed to align the strawberry with the end-effector. For early abort feedback during the deflating and snap-off stages, a micro-optical camera was embedded in the end-effector to detect and predict the probability of strawberry slippage in the gripper. This approach compensated for the absence of force feedback in the flexible pneumatic gripper, whose deformation during inflation and deflation rendered conventional force sensing impractical. However, a limitation of slip prediction–based early abort was that the material properties of artificial strawberries differed from those of real ones, particularly in weight and texture. Consequently, the secondary inflating-deflating regrasp experiment was evaluated only on the SnapData test set and was not conducted on HarvestFlex, owing to the material constraints of artificial strawberries. In this study, artificial strawberries were used primarily to evaluate relative error compensation and the early abort strategy. Mechanical failures, inverse kinematics errors, and other control-related faults were not addressed in this work. Due to the limitations of strawberry growth conditions, artificial strawberries were used in the experiments, which did not fully account for biological characteristics such as fruit damage.
Additionally, self-learning and adaptive evolutionary perception, planning and decision-making can be applied to reduce dataset dependency and improve the generalization and robustness of the strawberry-harvesting robot. With the rapid development of embodied artificial intelligence and robotic agents, a new wave of end-to-end, multi-modal, large-model-based perception, self-planning and self-decision frameworks is emerging in the robotics domain. These advances offer valuable insights for building strawberry harvesting robots and enabling collaborative operations among multiple robots. In the future, developing end-to-end active learning and continuously evolving methods is expected to become a major research trend.
@@ 6 Conclusion
This paper proposed a visual-based fault diagnosis and self-recovery system to address gripper offset and strawberry slippage during harvesting. An end-to-end multi-task perception network with a shared weight backbone and neck was developed to reduce computational overhead. Without the need for an additional external camera, relative error compensation method based on the simultaneous target-gripper detection was estimated, enabling compensation actions to mitigate cumulative errors. To monitor the strawberry’s status in the gripper, a miniature optical camera was embedded at its base. A MobileNet V3-Small classifier was adapted to detect strawberries within the gripper and trigger early abort signal when necessary during the deflating stage based on the GraspData. For strawberry slippage during the snap-off stage, a novel dataset, SnapData, comprising strawberry and background classes was introduced, and the perception network was extended and fine-tuned to segment strawberries, the gripper, and background. Width and height features were combined with a time-series LSTM classifier to predict strawberry slippage from the gripper, enabling early-abort picking actions and improving harvesting efficiency. Experimental results demonstrated that the perception network achieved high accuracy, with a mean absolute error of 0.035 for ripeness estimation. The compensation mechanism reduced relative errors to 3.12 mm and 4.11 mm, while the early-abort strategy shortened execution times by 1.89 s and 0.3 s during the deflating and snap-off stages, respectively. Overall, the proposed system effectively enhanced the robustness and efficiency of strawberry harvesting, providing a practical solution for reliable autonomous fruit picking.
@@
Data availability Data will be released.
@@
Code availability The code will be released.

### v2

@@ 5 Discussion
Vision-based fault diagnosis and self-recovery offered an effective means of enhancing the stability and efficiency of strawberry-harvesting robots. During the inflating and approaching stage, the positional relationship between the gripper and the harvesting point served as a visual indicator of accumulated errors caused by fruit recognition, localization, hand–eye calibration, and inverse kinematics. Without requiring an additional camera, relative errors based on the simultaneous target-gripper detection were computed to align the strawberry with the end-effector. For early abort feedback during the deflating and snap-off stages, a micro-optical camera was embedded in the end-effector to detect and predict the probability of strawberry slippage in the gripper. This approach compensated for the absence of force feedback in the flexible pneumatic gripper, whose deformation during inflation and deflation rendered conventional force sensing impractical. Mechanical failures, inverse kinematics errors, and other control-related faults were not addressed in this work.
Additionally, self-learning, continual learning and adaptive evolutionary perception, planning and decision-making can be applied to reduce dataset dependency and improve the generalization and robustness of the strawberry-harvesting robot. With the rapid development of embodied artificial intelligence and robotic agents, a new wave of end-to-end, multi-modal, large-model-based perception, self-planning and self-decision frameworks is emerging in the robotics domain. These advances offer valuable insights for building strawberry harvesting robots and enabling collaborative operations among multiple robots. In the future, developing end-to-end active learning and continuously evolving methods is expected to become a major research trend.
@@ 6 Conclusion
This paper proposed a visual-based fault diagnosis and self-recovery system to address gripper offset and strawberry slippage during harvesting. An end-to-end multi-task perception network with a shared weight backbone and neck was developed to reduce computational overhead. Without the need for an additional external camera, relative error compensation method based on the simultaneous target-gripper detection was estimated, enabling compensation actions to mitigate cumulative errors. To monitor the strawberry’s status in the gripper, a miniature optical camera was embedded at its base. A MobileNet V3-Small classifier was adapted to detect strawberries within the gripper and trigger early abort signal when necessary during the deflating stage based on the GraspData. For strawberry slippage during the snap-off stage, a novel dataset, SnapData, comprising strawberry and background classes was introduced, and the perception network was extended and fine-tuned to segment strawberries, the gripper, and background. Width, height and center point features were combined with a time-series LSTM classifier to predict strawberry slippage from the gripper on the SlipData, enabling early-abort picking actions and improving harvesting efficiency. Experimental results demonstrated that the perception network achieved high accuracy, with a mean absolute error of 0.035 for ripeness estimation. The compensation mechanism reduced relative errors to 3.12 mm and 4.11 mm, with a required execution time of less than 1 s. The response time of the grasp adjustment was approximately 0.4 s, resulting in a time saving of about 0.5 s during the deflating stage. For strawberry slipping status during the snap-off stage, re-inflation followed by a second snap-off required 1.81 $\pm$ 0.07 s and achieved a recovery success rate of 81.25%. When the operation was terminated due to slipped, the execution time was 1.44 $\pm$0.07 s, achieving a detection success rate of 88.89%. Compared with the original cycle time of 11.21 $\pm$ 0.15 s, grasp adjustment reduced the cycle time by approximately 4 s, and strawberry slip prediction reduced it by approximately 4 s as well. Overall, the proposed system effectively enhanced the robustness and efficiency of strawberry harvesting, providing a practical solution for reliable autonomous fruit picking. A demonstration video of visual diagnosis and self-recovery is available at https://1drv.ms/v/c/e68b602748988e28/IQAxE3rnaIYoTY9jHLkIULlnAZYotL9PTijSNry9i0aMnw4?e=HJhCL4.
@@
Data availability The data will be made publicly available.
@@
Code availability The code will be made publicly available.